Fermi National Laboratory
 

Accelerator Update for September 25, 2006 - September 27, 2006

 

Monday September 25

The day shift began with Operations monitoring a Recycler stash and the stack, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for colliding beam, with the Antiproton Source (Pbar) having stacking problems, and with an access into the Tevatron (TeV) sector holding off beam to the Linac, Booster, Main Injector (MI), and Recycler, NuMI and MiniBooNE.

At 8:16 AM, TeV RF experts reported finding a leak on TRF5. This will be repaired later.

Experts accessing into the TeV F sector reported at 8:22 AM that they found no cooling leaks on the kickers.

NuMI and MiniBooNE resumed taking beam at 9:23 AM.

At 9:52 AM, a TeV power supply (C:D0Q2) tripped on a fast bypass failure. Experts had the supply back on by 11:18 AM.

Operations put the TeV into shot setup at noon.

At 12:29 PM, Operation reported that the D0Q2 magnet quenched. Experts investigated and found a LCW leak in the power supply.

Operations reported at 2:27 PM that Pbar had resumed stacking.

Operations put the TeV through a dry squeeze at 3:46 PM.

Operations put the TeV into shot setup at 4:17 PM.

Operations established store 4975 at 7:56 PM with an initial luminosity of 186.8E30.

Tuesday September 26

The midnight shift began with Operations monitoring store 4975 (current average luminosity of 106.0 E30) and the stash (94.2E10), with Pbar stacking (13.4E10), and with NuMI and MiniBooNE taking beam.

At 1:18 AM, a MI power supply (I:SEXHPS) tripped off for the first of three times in an hour. It eventually reset each time.

At 2:06 AM, TRF7 alarmed.  Operators put its modulator in standby.

At 3:21 AM, I:SEXHPS tripped off and would not reset. Operators contacted experts. Pbar cannot stack and the Recycler cannot stash antiprotons without this power supply. A power supply expert reported at 4:59 AM that this supply had a ground fault. However, at 6:08 AM, experts managed to turn the supply on.

Pbar resumed stacking at 6:27 AM.

At 8:33 AM, store 4975 aborted; TRF1 had tripped off. There was no quench. Operations informed the RF experts.

Operations began preparing the TeV at 9:02 AM for an access.

D0 accessed their collision hall at 9:36 AM.

CDF accessed their collision hall at 10:08 AM.

Pbar experts began a stacktail study at 10:12 AM.

Recycler experts began a study at 10:13 AM.

Linac experts began work on KRF6 at 10:13 AM.

Operations began preparing the Booster at 10:17 AM for an access.

Operations prepared Pbar for an access at 11:34 AM.

A power supply expert began repairs on I:SEXHPS at 11:41 AM.

Booster came out of access at 1:11 PM.

Linac sent beam to the Booster at 2:21 PM and experts began a study period.

At 6:30 PM, Operations reported that I:SEXHPS wasn’t working.  Power supply experts were informed.

MiniBooNE resumed taking beam at 7:07 PM.

Experts finished their work on I:SEXHPS.  Operations began at 8:48 PM to establish beam to MI, NuMI, and Pbar.

Pbar resumed stacking at 9:07 PM.

At 10:57 PM, Operations reported that the F 17 kicker would not turn on due to a low cooling fluid indication. Operations contacted experts.

At 11:27 PM, Operations began preparing the TeV F sector for an access to check the kicker.

Wednesday September 27

The midnight shift began with Operations monitoring the stash, and with TeV experts preparing to access F sector.

At 1:55 AM, experts reported finding the cooling reservoir for F17KM2 dry; they filled the reservoir. There was a dripping leak on the upstream side of the magnet.

Operations resumed sending beam to Pbar, NuMI and MiniBooNE by 2:28 AM.

Operations put the TeV through a dry squeeze at 4:03 AM.

Operations put the TeV into shot setup at 5:31 AM.

At 7:50 AM, the TeV suffered an A1 and F4 quench during the ramp.

The Plans for Wednesday

The plans for today are to establish a store as soon as possible. F17KM2 needs to be replaced, which will take about 16 hours.  The Run Coordinator is planning for this work to begin on Thursday morning, 9/28/06. Other machines will be allowed to conduct work during this period.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/27/2006    email Fermilab